# M3GPT: An Advanced Multimodal, Multitask Framework for Motion Comprehension and Generation

### v4

@@ Abstract
This paper presents M3GPT, an advanced Multimodal, Multitask framework for Motion comprehension and generation. M3GPT operates on three fundamental principles. The first focuses on creating a unified representation space for various motion-relevant modalities. We employ discrete vector quantization for multimodal control and generation signals, such as text, music and motion/dance, enabling seamless integration into a large language model (LLM) with a single vocabulary. The second involves modeling model generation directly in the raw motion space. This strategy circumvents the information loss associated with discrete tokenizer, resulting in more detailed and comprehensive model generation. Third, M3GPT learns to model the connections and synergies among various motion-relevant tasks. Text, the most familiar and well-understood modality for LLMs, is utilized as a bridge to establish connections between different motion tasks, facilitating mutual reinforcement. To our knowledge, M3GPT is the first model capable of comprehending and generating motions based on multiple signals. Extensive experiments highlight M3GPT’s superior performance across various motion-relevant tasks and its powerful zero-shot generalization capabilities for extremely challenging tasks.
@@ 2 Related Work
Motion comprehension and Generation.  Motion comprehension involves two core tasks: motion-to-text and dance-to-music. Motion-to-text aims to describe human motion with natural language [31]. For example, recurrent networks have been used in [31] to accomplish this task. Dance-to-music involves creating a piece of music from a given dance [15, 21, 55]. For example, Zhun et al. [55] utilizes a generative adversarial network to generate music from dance videos. On the other hand, motion generation involves generating diverse human motions using multimodal inputs, such as text [37, 48, 13, 4, 49], music [21, 14, 44, 35] and incomplete motion [25, 1, 2]. Text-to-motion is one of the most important motion generation tasks. Recent works typically map text to motion using different architectures: diffusion model [49] and autoregressive transformer model [13]. Music-to-dance focuses on generating dance movements from music. For example, [35] predicts discrete token sequences conditioned on music, which are then used to regenerate the dance sequence. Motion Completion generates motion conditioning on partial motions, such as motion prediction [25, 1] and motion-in-between [28, 36]. Although these methods have shown promising results in various human motion tasks, most are limited to handling a single task. Until recently, some works [10, 16, 52, 53] attempt to integrate two or more tasks into a unified model, as shown in Tab. 3. However, these works either lack the ability of motion comprehension [53, 52] or fail to handle music modality [10, 16]. In this work, we propose a unified motion comprehension and generation framework that can handle multiple control signals simultaneously.
@@ 3.3 Training Strategy
  • Joint optimization of LLM and motion de-tokenizer.  Human motion (especially dance) encompasses intricate details. Previous works [16, 52] keep the motion de-tokenizer fixed during training LLM, which hinders LLM’s ability to perceive the distribution and details of motions. Specifically, in the output space of LLM, different motion tokens are treated as independent classes; therefore, the cost of classifying a motion token as semantic-similar token and semantic-distant token is the same. Apparently, relying solely on LLM’s autoregressive loss is insufficient for capturing the details of motion. To address this problem, we jointly optimize LLM and motion de-tokenizer in stage2 and stage3. This strategy enables the reconstruction error signals in raw motion space to backpropagate to LLM, enhancing LLM’s ability to generate the details of motion.
@@ 4 Experiments
4.1 Experimental setup
Implementation Details.  For motion tokenizer, we set the codebook size to $512$. As for music tokenizer, we use the pre-trained VQ-VAE from Jukebox with a codebook size of $2048$. In term of temporal downsampling rate, the motion encoder uses a rate of $4$, while the music encoder uses a rate of $128$. We utilize T5 base [33] as our language model backbone. For training the motion tokenizer, we use Adam as the optimizer with a batch size of $1000$ and an initial learning rate of $10^{-4}$. To train the language model backbone, we employ the Adafactor_dev optimizer and use CosineAnnealingLR as the learning rate scheduler. The learning rate is set to $2\times 10^{-4}$ for pre-training stage, and $10^{-4}$ for instruction fine-tuning stage. For hyperparameter settings, $\lambda$ in Eq. 4 is set to $0.2$, and $\beta$ in Eq. 2 is set to $0.02$. All our experiments are conducted on 8 NVIDIA A40 GPUs. Further details on implementation and hyperparameter analysis are provided in Appendix E.
@@ 4.1 Experimental setup
T2M-GPT [47] $\text{0.647}^{\pm\text{0.002}}$ $\text{0.101}^{\pm\text{0.005}}$ $\underline{\text{2.270}}^{\pm\text{0.033}}$
MotionDiffuse [49] $\underline{\text{0.659}}^{\pm\text{0.002}}$ $\textbf{\text{0.075}}^{\pm\text{0.004}}$ $\text{2.220}^{\pm\text{0.022}}$
@@
M3GPT (Instruction-tuned only T2M) $\textbf{\text{0.661}}^{\pm\text{0.003}}$ $\underline{\text{0.076}}^{\pm\text{0.002}}$ $\textbf{\text{2.273}}^{\pm\text{0.026}}$
@@
M3GPT (Pre-trained) $\underline{\text{27.65}}$ $\underline{\text{7.52}}$ 0.2105 92.35 7.67 0.2134 93.4 93.8 94.2
M3GPT (Instruction-tuned) 24.34 7.50 0.2056 86.47 7.75 0.2158 93.6 94.0 94.9
@@ 4.3 Comparisons with State-of-the-arts
Main results on text-to-motion dataset.  On the text-to-motion dataset, Motion-X, we evaluate M3GPT on $4$ tasks, i.e., text-to-motion, motion-to-text, motion prediction, and motion in-between. The comparison results are shown in Tab. 3. As shown, M3GPT achieves competitive performance across all evaluated tasks, highlighting its capability to address diverse motion tasks in a single model. Also, for text-to-motion task, M3GPT (instruction-tuned only T2M), which combines multitask pre-training and instruction fine-tuning solely on T2M task, yields better performance than Trained single task that only trains the model on T2M task. The results demonstrate that the multitask pre-training is conductive to enhance the performance of individual task.
Main results on music-to-dance datasets.  On the music-to-dance datasets, AIST++ and FineDance, we evaluate M3GPT on $2$ tasks, i.e., music-to-dance and dance-to-music. As shown in Tab. 4, in general, the performance of multitask pre-training and instruction fine-tuning in M3GPT outperforms single-task training, underscoring the effectiveness of multitask training for dance-relevant tasks. Also, M3GPT achieves competitive performance on most metrics. For music-to-dance, the best-performing method on the FineDance dataset is Lodge [20]. This approach features a specialized two-stage architecture for generating long-duration dance sequences, progressively refining movements from coarse to fine granularity using a diffusion model. On AIST++ dataset, M3GPT reports the best FIDk of $24.34$ for music-to-dance task, the best BHS and F1 of $94.0$ and $94.9$ for dance-to-music task.
@@ B.1 Data Introductions and Preprocessing
Motion-X includes 15.6 million precise 3D whole-body SMPL-X [30] pose annotations across 81.1K motion sequences with sequence-level semantic text descriptions. AIST++ contains 1,409 dance motion sequences across 10 dance genres with SMPL [24] pose annotations, and FineDance provides 7.7 hours of dance, totaling 831,600 frames with SMPL-H [34] poses at 30 fps across 16 dance genres.
@@ B.4 Tasks for Pre-training and Instruction tuning
Appendix C Quantitative Comparisons with SOTA Methods
In this section, we compare the performance of our M3GPT with existing SOTA methods on a broader set of metrics for each task across three datasets: Motion-X [23], AIST++ [18], and FineDance [19]. Tab. 9 shows the comparison of text-to-motion on Motion-X dataset. Tab. 9 shows the comparison of motion-to-text on Motion-X dataset. Tab. 9 shows the comparison of music-to-dance on AIST++ and FineDance datasets. Tab. 9 shows the comparison of dance-to-music on AIST++ and FineDance datasets. Tab. 10 shows the comparison of motion prediction and motion in-between on Motion-X dataset. As shown in these tables, M3GPT can achieve competitive performance with SOTAs across all evaluated tasks.
@@ Appendix C Quantitative Comparisons with SOTA Methods
Appendix D Qualitative Results
Fig. 6 presents visualizations for a variety of tasks, including text-to-motion, motion-to-text, motion prediction, motion in-between, music-to-dance, long-term dance generation, and music-text conditioned dance generation. The visualization results show that our method can generate realistic results across various motion-relevant tasks.
Appendix E Additional Experiments

### v5

@@ Abstract
This paper presents M3GPT, an advanced Multimodal, Multitask framework for Motion comprehension and generation. M3GPT operates on three fundamental principles. The first focuses on creating a unified representation space for various motion-relevant modalities. We employ discrete vector quantization for multimodal conditional signals, such as text, music and motion/dance, enabling seamless integration into a large language model (LLM) with a single vocabulary. The second involves modeling motion generation directly in the raw motion space. This strategy circumvents the information loss associated with a discrete tokenizer, resulting in more detailed and comprehensive motion generation. Third, M3GPT learns to model the connections and synergies among various motion-relevant tasks. Text, the most familiar and well-understood modality for LLMs, is utilized as a bridge to establish connections between different motion tasks, facilitating mutual reinforcement. To our knowledge, M3GPT is the first model capable of comprehending and generating motions based on multiple signals. Extensive experiments highlight M3GPT’s superior performance across various motion-relevant tasks and its powerful zero-shot generalization capabilities for extremely challenging tasks. Project page: https://github.com/luomingshuang/M3GPT.
@@ 2 Related Work
Motion comprehension and Generation.  Many existing works focus on studying human appearance, pose, detection, attribute, part parsing and so on [61, 19, 45, 40, 23, 17]. This work focuses on studying human motion, including motion comprehension and motion generation. Motion comprehension involves two core tasks: motion-to-text and dance-to-music. Motion-to-text aims to describe human motion with natural language [37]. For example, recurrent networks have been used in [37] to accomplish this task. Dance-to-music involves creating a piece of music from a given dance [20, 27, 64]. For example, Zhun et al. [64] utilizes a generative adversarial network to generate music from dance videos. On the other hand, motion generation involves generating diverse human motions using multimodal inputs, such as text [44, 56, 15, 5, 57], music [27, 18, 52, 42] and incomplete motion [31, 1, 3]. Text-to-motion is one of the most important motion generation tasks. Recent works typically map text to motion using different architectures: diffusion model [57] and autoregressive transformer model [15]. Music-to-dance focuses on generating dance movements from music. For example, [42] predicts discrete token sequences conditioned on music, which are then used to regenerate the dance sequence. Motion Completion generates motion conditioning on partial motions, such as motion prediction [31, 1] and motion-in-between [34, 43]. Although these methods have shown promising results in various human motion tasks, most are limited to handling a single task. Until recently, some works [12, 21, 60, 62] attempt to integrate two or more tasks into a unified model, as shown in Tab. 3. However, these works either lack the ability of motion comprehension [62, 60] or fail to handle music modality [12, 21]. In this work, we propose a unified motion comprehension and generation framework that can handle multiple control signals simultaneously.
@@ 3.3 Training Strategy
  • Joint optimization of LLM and motion de-tokenizer.  Human motion (especially dance) encompasses intricate details. Previous works [21, 60] keep the motion de-tokenizer fixed during training LLM, which hinders LLM’s ability to perceive the distribution and details of motions. Specifically, in the output space of LLM, different motion tokens are treated as independent classes; therefore, the cost of classifying a motion token as semantic-similar token and semantic-distant token is the same. Apparently, relying solely on LLM’s autoregressive loss is insufficient for capturing the details of motion. To address this problem, we jointly optimize LLM and motion de-tokenizer in stage2 and stage3. This strategy enables the reconstruction error signals in raw motion space to backpropagate to LLM, enhancing LLM’s ability to generate the details of motion. With the goal of minimizing L1 loss between the predicted and real motion, we search for the motion’s token sequence that could minimize this L1 loss in original motion space. As the motion de-tokenizer continuously optimizes, the target motion’s token sequence, which supervises LLM training, dynamically changes. This dynamic adjustment reduces L1 loss progressively, achieving joint optimization.
@@ 4 Experiments
4.1 Experimental setup
Implementation Details.  For motion tokenizer, we set the codebook size to $512$. As for music tokenizer, we use the pre-trained VQ-VAE from Jukebox with a codebook size of $2048$. In term of temporal downsampling rate, the motion encoder uses a rate of $4$, while the music encoder uses a rate of $128$. We utilize T5 base [39] as our language model backbone. For training the motion tokenizer, we use Adam as the optimizer with a batch size of $1000$ and an initial learning rate of $10^{-4}$. To train the language model backbone, we employ the Adafactor_dev optimizer and use CosineAnnealingLR as the learning rate scheduler. The learning rate is set to $2\times 10^{-4}$ for pre-training stage, and $10^{-4}$ for instruction fine-tuning stage. For hyperparameter settings, $\lambda$ in Eq. 4 is set to $0.2$, and $\beta$ in Eq. 2 is set to $0.02$. All our experiments are conducted on 8 NVIDIA A40 GPUs. To evaluate the model’s performance across different platforms, we also test our trained M3GPT with T5-base on Ascend 910B NPUs. Further details on implementation and hyperparameter analysis are provided in Appendix E.
@@ 4.1 Experimental setup
T2M-GPT [55] $\text{0.647}^{\pm\text{0.002}}$ $\text{0.101}^{\pm\text{0.005}}$ $\underline{\text{2.270}}^{\pm\text{0.033}}$ 0.814 1.755 - -
MotionDiffuse [57] $\text{0.659}^{\pm\text{0.002}}$ $\underline{\text{0.075}}^{\pm\text{0.004}}$ $\text{2.220}^{\pm\text{0.022}}$
@@
MoMask [16] $\textbf{\text{0.668}}^{\pm\text{0.003}}$ $\textbf{\text{0.074}}^{\pm\text{0.004}}$ $\text{2.241}^{\pm\text{0.016}}$ 0.626 1.884
MotionLCM [8] $\text{0.658}^{\pm\text{0.005}}$ $\text{0.078}^{\pm\text{0.003}}$ $\text{2.206}^{\pm\text{0.026}}$
MotionGPT[21] $\text{0.659}^{\pm\text{0.003}}$ $\text{0.078}^{\pm\text{0.001}}$ $\text{2.166}^{\pm\text{0.026}}$ $\underline{\text{0.840}}$ 11.21 31.36 0.701 1.818 0.648 1.875
@@
M3GPT (Instruction-tuned only T2M) $\underline{\text{0.661}}^{\pm\text{0.003}}$ $\text{0.076}^{\pm\text{0.002}}$ $\textbf{\text{2.273}}^{\pm\text{0.026}}$
Table 4: Comparison results on Motion-X [29] dataset based on Ascend 910B NPUs.
[…]
Trained single task $\underline{\text{0.654}}^{\pm\text{0.002}}$ $\underline{\text{0.081}}^{\pm\text{0.003}}$ $\underline{2.304}^{\pm\text{0.017}}$ 0.763 10.16 22.89 0.776 1.818 0.712 1.880
M3GPT (Pre-trained) $\text{0.596}^{\pm\text{0.002}}$ $\text{0.096}^{\pm\text{0.003}}$ $\text{2.241}^{\pm\text{0.018}}$ 0.831 11.05 24.03 0.710 1.882 0.608 1.874
M3GPT (Instruction-tuned) $\text{\text{0.612}}^{\pm\text{0.002}}$ $\text{0.094}^{\pm\text{0.002}}$ $\text{2.276}^{\pm\text{0.021}}$ 0.846 11.52 42.64 0.684 1.841 0.621 1.903
M3GPT (Instruction-tuned only T2M) $\textbf{\text{0.659}}^{\pm\text{0.003}}$ $\textbf{\text{0.078}}^{\pm\text{0.003}}$ $\textbf{\text{2.314}}^{\pm\text{0.023}}$
@@
M3GPT (Pre-trained) 27.65 7.52 0.2105 92.35 7.67 0.2134 93.4 93.8 94.2
M3GPT (Instruction-tuned) $\underline{\text{24.34}}$ 7.50 0.2056 86.47 7.75 0.2158 93.6 94.0 94.9
M3GPT (Instruction-tuned for single task) 23.01 7.85 0.2261 42.66 8.24 0.2231 94.3 94.0 95.0
Table 6: Comparison results on AIST++ [24] and FineDance [25] datasets based on Ascend 910B NPUs.
[…]
Trained single task 77.32 5.61 0.1860 134.66 6.52 0.2088 93.8 93.6 92.7
M3GPT (Pre-trained) 27.99 $\underline{\text{7.61}}$ $\underline{\text{0.2102}}$ 91.39 7.71 0.2128 93.4 93.6 94.2
M3GPT (Instruction-tuned) $\underline{\text{25.05}}$ 7.48 0.2072 88.25 7.76 0.2160 93.5 93.8 94.7
M3GPT (Instruction-tuned for single task) 23.68 7.83 0.2264 43.78 8.39 0.2225 94.3 94.0 94.9
@@ 4.3 Comparisons with State-of-the-arts
Main results on text-to-motion dataset.  On the text-to-motion dataset, Motion-X, we evaluate M3GPT on $4$ tasks, i.e., text-to-motion, motion-to-text, motion prediction, and motion in-between. The comparison results are shown in Tab. 6. As shown, M3GPT achieves competitive performance across all evaluated tasks, highlighting its capability to address diverse motion tasks in a single model. Also, for text-to-motion task, M3GPT (instruction-tuned only T2M), which combines multitask pre-training and instruction fine-tuning solely on T2M task, yields better performance than Trained single task that only trains the model on T2M task. Tab. 6 presents the results of testing on the Ascend 910B NPUs, where M3GPT also achieves comparably good performance. These results demonstrate that multitask pre-training can enhance the performance of individual tasks across different computation platforms. Further results of M3GPT trained on NPUs will be presented later.
Main results on music-to-dance datasets.  On the music-to-dance datasets, AIST++ and FineDance, we evaluate M3GPT on $2$ tasks, i.e., music-to-dance and dance-to-music. As shown in Tab. 6, in general, the performance of multitask pre-training and instruction fine-tuning in M3GPT outperforms single-task training, underscoring the effectiveness of multitask training for dance-relevant tasks. Also, M3GPT achieves competitive performance on most metrics. For music-to-dance, the best-performing method on the FineDance dataset is Lodge [26]. This approach features a specialized two-stage architecture for generating long-duration dance sequences, progressively refining movements from coarse to fine granularity using a diffusion model. On AIST++ dataset, M3GPT reports the best FIDk of $24.34$ for music-to-dance task, the best BHS and F1 of $94.0$ and $94.9$ for dance-to-music task. The results in Tab. 6, tested on the Ascend 910B NPUs, also demonstrate that multitask training can enhance the performance of both music-to-dance and dance-to-music tasks.
@@ 5 Conclusion
Acknowledgments. This work is sponsored by the National Natural Science Foundation of China (NSFC) under Grant 62306301, the National Postdoctoral Program for Innovative Talents under Grant BX20220310, and the CAAI-CANN Open Fund developed on the OpenI Community. It is also supported by the project of Peng Cheng Laboratory (PCL2023A08).
@@ B.1 Data Introductions and Preprocessing
Motion-X includes 15.6 million precise 3D whole-body SMPL-X [36] pose annotations across 81.1K motion sequences with sequence-level semantic text descriptions. AIST++ contains 1,409 dance motion sequences across 10 dance genres with SMPL [30] pose annotations, and FineDance provides 7.7 hours of dance, totaling 831,600 frames with SMPL-H [41] poses at 30 fps across 16 dance genres. Tab. 7 shows the training datasets and their corresponding sample numbers that we use to train our model.
@@
Table 7: The training datasets and sample numbers for different tasks.
Tasks T2M M2T Motion Prediction/In-between A2D D2A A2T T2D
Training dataset Motion-X Motion-X Motion-X/AIST++/FineDance AIST++/FineDance AIST++/FineDance AIST++/FineDance AIST++/FineDance
Training samples number 64867 64867 64867/952/177 952/177 952/177 952/177 952/177
@@ B.4 Tasks for Pre-training and Instruction tuning
Appendix C Quantitative Results and Comparisons with SOTA Methods
In this section, we compare the performance of our M3GPT with existing SOTA methods on a broader set of metrics for each task across three datasets: Motion-X [29], AIST++ [24], and FineDance [25]. Tab. 12 shows the comparison of text-to-motion on Motion-X dataset. Tab. 12 shows the comparison of motion-to-text on Motion-X dataset. Tab. 12 shows the comparison of music-to-dance on AIST++ and FineDance datasets. Tab. 12 shows the comparison of dance-to-music on AIST++ and FineDance datasets. Tab. 18 and Tab. 18 shows the comparison of motion prediction and motion in-between on Motion-X dataset. Tab. 18 shows the comparison of music-to-text, text-to-dance and text-to-music on AIST++ dataset. Tab. 18 and Tab. 18 show the comparison of motion-related tasks among different size of T5. As shown in these tables, M3GPT can achieve competitive performance with SOTAs across all evaluated tasks.
Appendix D Qualitative Results and Comparison with SOTA Methods
Fig. 6 presents visualizations for a variety of tasks, including text-to-motion, motion-to-text, motion prediction, motion in-between, music-to-dance, long-term dance generation, and music-text conditioned dance generation. The visualization results show that our method can generate realistic results across various motion-relevant tasks. Fig. 7 presents the qualitative results between different methods for text-to-motion and music-to-dance.
Appendix E Additional Experiments
@@
Table 14: Comparison of MPJPE on Motion Prediction and Motion In-between on Motion-X [29].
Methods on Motion-X Motion Prediction (MPJPE $\downarrow$) Motion In-between (MPJPE $\downarrow$)
T2M-GPT [55] 80.2 63.7
MoMask [16] 67.9 55.2
MotionGPT [21] 71.3 59.9
M3GPT (Instruction-tuned) 54.2 51.0
Table 15: Comparison of Music-to-Text (A2T), Text-to-Dance (T2D) and Text-to-Music (T2A) on AIST++.
Methods on AIST++ Music-to-Text Text-to-Dance Text-to-Music
Bleu@4 $\uparrow$ CIDEr $\uparrow$ R-TOP1 $\uparrow$ FID $\downarrow$ BCS $\uparrow$ BHS $\uparrow$
M3GPT (Single task training for A2T) 9.24 24.55 - - - -
M3GPT (Single task training for T2D) - - 0.541 0.095 - -
MusicLDM [2] - - - - 74.5 73.8
Mubert - - - - 73.3 73.0
M3GPT (Instruction-tuned) 11.95 28.88 0.588 0.077 74.5 74.7
Table 16: Comparison of Text-to-Motion and Motion-to-Text with different size of T5.
Methods on Motion-X LLM Training time Text-to-Motion Motion-to-Text
R-TOP1 $\uparrow$ FID $\downarrow$ Div $\uparrow$ R-TOP3 $\uparrow$ Bleu4 $\uparrow$ CIDEr $\uparrow$
M3GPT T5-small (60M) 5 days 0.598 0.096 2.202 0.822 10.43 38.22
M3GPT T5-base (220M) 7 days 0.615 0.093 2.253 0.845 11.50 42.93
M3GPT T5-large (770M) 8 days 0.619 0.090 2.256 0.848 11.64 43.05
Table 17: Comparison of Music-to-Dance and Dance-to-Music with different size of T5.
Methods on AIST++ LLM Training time Music-to-Dance Dance-to-Music
FIDk $\downarrow$ DIVk $\uparrow$ BAS $\uparrow$ BCS $\uparrow$ BHS $\uparrow$
M3GPT T5-small (60M) 5 days 28.05 5.96 0.2091 89.1 91.2
M3GPT T5-base (220M) 7 days 24.34 7.50 0.2056 93.6 94.0
M3GPT T5-large (770M) 8 days 23.26 7.54 0.2061 93.8 94.1